3. b. Calculate the elasticity of supply if a 15 %increase in the price of soya bean oil increases its supply from 300 to 345 units (5 Marks) 



1
Expert's answer
2021-04-27T07:28:36-0400
"E_s=\dfrac{\%\Delta Q}{\%\Delta P}=\dfrac{\dfrac{Q_2-Q_1}{0.5(Q_2+Q_1)}}{\%\Delta P},""E_s=\dfrac{\%\Delta Q}{\%\Delta P}=\dfrac{\dfrac{345-300}{0.5(345+300)}}{0.15}=0.93"